A mixture contains 5.432 g of sand and 2.521 g of salt. What is the percentage of salt in the mixture?
Answer #1

Sol . As Mass of sand = 5.432 g

and Mass of salt = 2.521 g

So , Total mass of mixture = Mass of sand + Mass of salt = 5.432 g + 2.521 g = 7.953 g

Therefore , Percentage of salt in the mixture

= (Mass of salt × 100)/Total mass of mixture

= (2.521 × 100)/7.953 =  31.70 %